# Depth and slip ratio dependencies of friction for a sphere rolling on a granular slope

### v1

@@ 1 Introduction
For the case of pure translational motion relating to drag force and the resistance of granular bed, many investigations were reported [14, 19, 23, 2, 11, 9, 21, 25, 24, 20, 17, 18, 3, 4]. For instance, the drag force during vertical impact cratering into granular media was analyzed [14, 19]. Penetration drag at constant velocity was also examined [23, 2]. It was shown that, depending on the object’s shape, the resistance was proportional to either the square or the cube of its radius [2]. Iikawa and Katsuragi explored the effects of particle adhesion (wet granular), inter-particle friction, and shape of intrusion on penetration resistance [11]. Objects moving at constant velocity through granular media were also examined [9, 21, 25], where it was found that horizontally moving objects experienced both horizontal resistance and vertical lift forces. Motions with acceleration [24] or deceleration [20] in granular media were also considered. Translational motion accompanied by a certain degree of sinking into the granular surface, either at an imposed constant velocity [17, 18, 3] or under gravity along granular slopes [4], was likewise reported. These studies demonstrated that deformation of the granular medium affected the friction between the object and the surface [17, 18, 4].
@@
In the case of combination of rotating and translating motion, some studies were conducted [13, 26, 29, 8, 6, 10]. Penetration with rotation at constant velocity was also investigated [13], where it was found that rotational motion reduces the drag force. Impact cratering with rolling was likewise examined [13], and the results showed that rotation increases the sinking depth. Motion with rolling at various slip ratios was explored [26], demonstrating that rotational motion reduced the drag force up to a certain rotational velocity. The characteristic of these studies [7, 13, 26] is that they are externally driven and objects move at a constant speed. Some studies addressed the behavior of a sphere rolling on granular surface without its translational or rotational speed being controlled by a motor. In this perspective, cases such as rolling on a flat horizontal granular surface [29], rolling down a granular slope with deceleration [8], rolling down with acceleration [6], and rolling up a granular slope [10] are of direct relevance to.
@@
Wal et al. examined the friction between a flat horizontal granular surface and a rolling basketball or medicine ball using both numerical and experimental approaches [29]. Although they suggested that the friction coefficient depends on the sinking depth, no quantitative relationship was obtained. Fukumoto et al. analyzed the dynamics of a sphere rolling up a granular slope [10] and found an empirical quantitative relation in which the friction coefficient is proportional to the sinking depth. However, the range of applicability of this empirical relation, namely its dependence on the slope angle (including both rolling up and down), was not clarified. In fact, the friction coefficient observed when a sphere rolled down a granular slope [8] was larger than the value expected from this empirical law. Notably, Blasio and Saeter [8] conducted their experiments only within a narrow range of sphere densities (glass and steel) and slope angles (14.5–20.5). Thus, the dependencies on density and slope angle were not fully clarified.
@@
In this study, we experimentally investigate the effect of a wide range of sphere densities and slope angles on the decelerative dynamics and friction of a sphere rolling down a granular slope. The friction is characterized from kinematic data obtained in a dedicated experimental setup. The details of the experimental apparatus, parameters, and conditions are described in Sec. 2. The experimental results analyzed using the analytical method outlined in Sec. 3 are presented in Sec. 4. A discussion introducing the frictional framework is given in Sec. 5, and the Conclusion is provided in Sec. 6
@@ 2 Experiment
To obtain the dynamics of the sphere rolling down a granular slope, we build the experimental setup as shown in Fig. 1 (a). The sphere is released at various heights on the aluminium rail (rail width: 6 $\mathrm{mm}$) to obtain the initial velocity $v_{0}$ (0.3–0.9 $\mathrm{m/s}$ as shown in tab. 1) at $X=0$, and then enters the granular slope. The $X$ axis is taken along the granular slope. The spheres used in this experiment are made of polyethylene, polyacetal, glass, and alumina ceramic. Their densities are $\rho_{\mathrm{s}}=930$, 1400, 2600, and 3900 $\mathrm{kg/m^{3}}$, respectively. The spheres used in this study are the same as those used in the previous uphill study [10]. However, the upper density limit is set to that of alumina ceramic since it was impossible to perform image analysis for a stainless steel sphere due to the large splash associated with its deep sinking [10]. All these spheres have an identical radius of $R=6.35$ mm. The box (width: 200 mm, length: 100 mm, height: 50 mm) filled with glass beads is inclined by an angle $\alpha$. The $\alpha$ is varied as $\alpha\simeq-5^{\circ}$, $10^{\circ}$, $15^{\circ}$, or $20^{\circ}$. The typical diameter of the glass beads used in this experiment is 0.8 mm (770–910 µm, AS-One, BZ-08). The true density of the glass beads is $\rho=2500~\mathrm{kg/m^{3}}$. The bulk density of the granular layer in this experiment is estimated as $\rho_{\mathrm{g}}=1600\pm 40~\mathrm{kg/m^{3}}$, which corresponds to the packing fraction $\phi$ given by $\rho_{\mathrm{g}}=\phi\rho$, resulting in $\phi=0.64\pm 0.02$. Each experimental condition has five experimental runs to evaluate the reproducibility. Before each run, the granular layer is completely removed from the box and the box is refilled with glass beads. To capture the dynamics of the sphere, we use a high-speed camera (Omron Sentech, STCMCCM401U3V) operating at 200 fps with a spatial resolution of 0.13 mm/pixel (image size: 2048 × 700 pixels). From the acquired images, we measure the instantaneous position and rolling posture of the sphere, the sinking depth $\delta$, the maximum travel distance $L$, and the slip ratio $s$ ($s=1-L/R\theta_{\mathrm{stop}}$, where $\theta_{\mathrm{stop}}$ is the total rotational angle) as shown in Fig. 1 (b).
@@
Figure 1: (a) Schematic view of the experimental setup. By rolling down the rail made of aluminium, the sphere of the radius $R$ obtains initial translational velocity $v_{0}$. The rolling posture $\theta$ is measured counterclockwise. The sphere rolls down the granular slope of the angle $\alpha$. $\alpha<0$ and $\alpha>0$ (obtained by Fukumoto et al. [10]) are defined as downhill and uphill, respectively. (b) The sphere moves in the $X$ direction by a distance $L$ and sinks by a depth $\delta$ when it stops. [(a), (b)] The size ratio between the sphere and the glass beads is not to scale.
@@
$\alpha=-5^{\circ}$ 0.32-0.55 0.32-0.42 0.35-0.40 0.35-0.45
$\alpha=-10^{\circ}$ 0.45-0.7 0.45-0.78 0.56-0.9 0.6-0.95
$\alpha=-15^{\circ}$ 0.37-0.60 0.36-0.67 0.46-0.7 0.6-0.85
$\alpha=-20^{\circ}$ 0.35-0.5 0.43-0.6 0.44-0.7 0.5-0.77
@@ 3 Analysis
The typical examples of actual movies are provided as supplementary material. As shown in Fig. 2 (a), a hemisphere of each sphere, except for the glass sphere, is painted red to detect the rolling orientation. Due to the surface condition of the glass sphere, it is not possible to paint it red with a permanent marker. The rotational angle, $\theta_{\mathrm{stop}}$, before stopping is analyzed in the counterclockwise direction by detecting the slope of the boundary between the white and red regions except for the glass sphere. By detecting the center of the sphere in each image, as shown in Fig. 2 (b), we obtain the sinking depth $\delta$ in the direction of gravity and the translational trajectory of the sphere, $X(t)$, where $t$ is the elapsed time after entering the granular slope.
@@ 4 Result
The measured sinking depth normalized by the sphere radius, $\delta/R$, is plotted in Fig. 3. As shown in Fig. 3 (a), $\delta/R$ is almost independent of $v_{0}$. Therefore, in Fig. 3 (b)–(e), the average is taken over data with different $v_{0}$, and it can be seen that it does not depend on $\alpha$, similar to $v_{0}$. However, the average value of $\delta/R$ depends on the density of the sphere, $\rho_{\mathrm{s}}$. The density dependence of sinking depth has been investigated previously [28, 6, 10]. For a sphere dropped onto a granular surface without initial velocity, the normalized sinking depth was found to scale with the density ratio $(\rho_{\mathrm{s}}/\rho_{\mathrm{g}})$ as obtained by Uehara et al. [28],
@@
where $C_{\rho}$ is a dimensionless constant that depends on the friction coefficient between grains. Uehara et al. reported $C_{\rho}=0.51$ for a flat horizontal glass-bead surface [28], while Texier et al. obtained $C_{\rho}=0.61$ for a granular slope [6]. Although the value of $C_{\rho}$ differs, this scaling law [Eq. (1)] can also be applied to the sinking depth after a sphere rolled up a granular slope ($C_{\rho}=0.46\pm 0.05$[10]. When applying the same scaling to the rolling-down case, we find that the relation is satisfied with $C_{\rho}=0.38\pm 0.06$, as shown in Fig. 4. Therefore, this scaling law is valid for spheres rolling on granular surfaces, and also the coefficient $C_{\rho}$ does not vary significantly between the rolling-down and rolling-up cases [10].
@@
Figure 4: The double logarithmic plot of $\delta/R$ vs $\rho_{\mathrm{s}}/\rho_{\mathrm{g}}$ for both downhill and uphill. We use the data (uphill) obtained by Fukumoto et al. [10]. Error bars indicate the standard deviation of various $\alpha,v_{0}$ conditions. The blue and red dashed line indicate the scaling relation [Eq. (1)] with $C_{\mathrm{\rho}}=0.38$ and 0.46, respectively.
@@
The slip ratio $s$ is computed by substituting the measured values of $L$ and $\theta_{\mathrm{stop}}$ into Eq. (2). Note that when the sphere is rolling down a slope, the translational and rotational motions stop almost simultaneously. As shown in Fig. 7, $s$ increases with $\alpha$, where the definition of $\pm\alpha$ corresponds to rolling up $(+)$ and rolling down ()(-). (For rolling up $(+)$, we use the data obtained by Fukumoto et al. [10].) Moreover, the variation of $s$ appears to be continuous for each sphere as $\alpha$ changes.
As shown in Fig. 8, a bump is formed in front of the sphere, and we will focus on the relationship between this bump and slip ratio. Even though the sinking depth changes little between rolling up and rolling down, the deformation of the granular surface in front of the sphere is larger when it rolls down a granular slope, as shown in Fig. 8. During rolling down motion, the sphere is supposed to accelerate under gravity but the sphere actually decelerates due to the effect of the bump of glass beads accumulating in front of it. The bumpy structure produces a resistive force opposing its motion. In contrast, during rolling up motion, the deceleration due to gravity dominates over the resistance from the granular bed. These observations indicate that the size of bump is larger during rolling down motion than during rolling up motion. Numerical calculations have shown that the larger the slip ratio of a wheel rolling on granular ground, the smaller the bump formation (terrain deformation formed in front of the wheel) [15]. This tendency between bump formation and slip ratio is consistent with our observations.
Figure 7: The relation between slip ratio $s$ and angle $\alpha$. $\alpha<0$ and $\alpha>0$ correspond to rolling down and up cases respectively. We use the data ($\alpha>0$) obtained by Fukumoto et al. [10]. Error bars indicate the standard deviation of all the data.

### v2

@@ 1 Introduction
For the case of pure translational motion relating to drag force and the resistance of granular bed, many investigations were reported. For instance, the drag force during vertical impact cratering into granular media was analyzed [14, 20]. Penetration drag at constant velocity was also examined [24, 2]. It was shown that, depending on the object’s shape, the resistance was proportional to either the square or the cube of its radius [2]. Iikawa and Katsuragi explored the effects of particle adhesion (wet granular), inter-particle friction, and shape of intrusion on penetration resistance [11]. Objects moving at constant velocity through granular media were also examined [9, 22, 26], where it was found that horizontally moving objects experienced both horizontal resistance and vertical lift forces. Motions with acceleration [25] or deceleration [21] in granular media were also considered. Translational motion accompanied by a certain degree of sinking into the granular surface, either at an imposed constant velocity [17, 18, 3] or under gravity along granular slopes [4], was likewise reported. These studies demonstrated that deformation of the granular medium affected the friction between the object and the surface [17, 18, 4].
@@
In the case of combination of rotating and translating motion, some studies were conducted. Penetration with rotation at constant velocity was also investigated [13], where it was found that rotational motion reduces the drag force. Impact cratering with rolling was likewise examined [13], and the results showed that rotation increases the sinking depth. Motion with rolling at various slip ratios was explored [27], demonstrating that rotational motion reduced the drag force up to a certain rotational velocity. The characteristic of these studies [7, 13, 27] is that they are externally driven and objects move at a constant speed. Some studies addressed the behavior of a sphere rolling on granular surface without its translational or rotational speed being controlled by a motor. In this perspective, cases such as rolling on a flat horizontal granular surface [30], rolling down a granular slope with deceleration [8], rolling down with acceleration [6], and rolling up a granular slope [10] are of direct relevance to the present study.
@@
Wal et al. examined the friction between a flat horizontal granular surface and a rolling basketball or medicine ball using both numerical and experimental approaches [30]. Although they suggested that the friction coefficient depends on the sinking depth, no quantitative relationship was obtained. Fukumoto et al. analyzed the dynamics of a sphere rolling up a granular slope [10] and found an empirical quantitative relation in which the friction coefficient $\mu_{\mathrm{d}}$ is proportional to the sinking depth normalized to the radius of the sphere $\delta/R$, $\mu_{\mathrm{d}}=0.49(\delta/R)$. However, the range of applicability of this empirical relation, namely its dependence on the slope angle (including both rolling up and down), was not clarified. In fact, the friction coefficient observed when a sphere rolled down a granular slope [8], $\mu_{\mathrm{d}}\simeq 0.45$ was larger than the value expected from this empirical law, $\mu_{\mathrm{d}}=0.49(\delta/R)$, because the normalized sinking depth was $\delta/R\simeq 0.3$. Besides, Blasio and Saeter [8] conducted their experiments only within a narrow range of sphere densities (they only use glass and steel) and slope angles (14.5–20.5). Thus, the dependencies on density and slope angle were not fully clarified.
@@
In this study, we experimentally investigate the effect of a wide range of sphere densities and slope angles on the decelerative dynamics and friction of a sphere rolling down a granular slope. In other words, applicable range of the previously reported empirical law is experimentally evaluated. The results suggest that the normalized depth and slip ratio determine the effective friction of a rolling sphere. The friction is characterized from kinematic data obtained in a dedicated experimental setup. The details of the experimental apparatus, parameters, and conditions are described in Sec. 2. The experimental results analyzed using the analytical method outlined in Sec. 3 are presented in Sec. 4. A discussion introducing the frictional framework is given in Sec. 5, and the Conclusion is provided in Sec. 6
@@ 2 Experiment
To obtain the dynamics of the sphere rolling down a granular slope, we build the experimental setup as shown in Fig. 1 (a) and (b). The sphere is released at various heights on the aluminium rail (rail width: 6 $\mathrm{mm}$) to obtain the initial velocity $v_{0}$ (0.32–0.95 $\mathrm{m/s}$ as shown in tab. 1) at $X=0$, and then enters the granular slope. The initial velocity $v_{0}$ is varied by changing the initial release position of the sphere. We intentionally adjust $v_{0}$ so that the sphere would come to rest within the glass-bead box mentioned later in this section. In addition, the minimum initial velocity is limited such that the sphere rolls over a sufficient distance (>> 20 mm) to minimize wall effects. Because the deceleration depends on the density of the sphere, it is necessary to adjust the initial velocity $v_{0}$ to some extent. The $X$ axis is taken along the granular slope. The spheres used in this experiment are made of polyethylene, polyacetal, glass, and alumina ceramic (mainly supplied by As-one). Their densities are $\rho_{\mathrm{s}}=930$, 1400, 2600, and 3900 $\mathrm{kg/m^{3}}$, respectively. All spheres are sufficiently rigid to neglect deformation and have the identical surface static friction coefficients against the fixed glass beads surface, 0.12. This value is measured by the critical angle on which the sphere start to roll down on the fixed glass beads surface. The spheres used in this study are the same as those used in the previous uphill study [10]. However, the upper density limit is set to that of alumina ceramic since it was impossible to perform image analysis for a stainless steel sphere due to the large splash associated with its deep sinking [10]. The lower density limit corresponds to that of polyethylene sphere because the lighter density (e.g. polystyrene) sphere cannnot be ignored in the air resistance. The density range is determined based on these criteria. All these spheres have an identical radius of $R=6.35$ mm. The spheres used in the experiments have diameters approximately ten times larger than the particle size, ensuring a scale much larger than that of individual grains. The box (width: 200 mm, length: 100 mm, height: 50 mm) filled with glass beads is inclined by an angle $\alpha$. The $\alpha$ is varied as $\alpha\simeq-5^{\circ}$, $10^{\circ}$, $15^{\circ}$, or $20^{\circ}$. The typical diameter of the glass beads used in this experiment is 0.8 mm (770–910 µm, AS-One, BZ-08). The employed diameter should be large enough to neglect air drag and cohesive forces. At the same time, the size ratio between the rolling sphere and the glass beads should be large. Considering this compromise, the sizes of the materials used were determined. The true density of the glass beads is $\rho=2500~\mathrm{kg/m^{3}}$. The bulk density of the granular layer in this experiment is estimated as $\rho_{\mathrm{g}}=1600\pm 40~\mathrm{kg/m^{3}}$, which corresponds to the packing fraction $\phi$ given by $\rho_{\mathrm{g}}=\phi\rho$, resulting in $\phi=0.64\pm 0.02$. Each experimental condition has five experimental runs to evaluate the reproducibility. Before each run, the granular layer is completely removed from the box and the box is refilled with glass beads. To capture the dynamics of the sphere, we use a high-speed camera (Omron Sentech, STCMCCM401U3V) operating at 200 fps with a spatial resolution of 0.13 mm/pixel (image size: 2048 × 700 pixels). From the acquired images, we measure the instantaneous position and rolling posture of the sphere, the sinking depth $\delta$, the maximum travel distance $L$, and the slip ratio $s$ ($s=1-L/R\theta_{\mathrm{stop}}$, where $\theta_{\mathrm{stop}}$ is the total rotational angle) as shown in Fig. 1 (c).
@@
Figure 1: (a) Schematic view of the experimental setup. By rolling down the rail made of aluminium, the sphere of the radius $R$ obtains initial translational velocity $v_{0}$. The rolling posture $\theta$ is measured counterclockwise. The sphere rolls down the granular slope of the angle $\alpha$. $\alpha<0$ and $\alpha>0$ (obtained by Fukumoto et al. [10]) are defined as downhill and uphill, respectively. (b) The view of the experimental setup. (c) The sphere moves in the $X$ direction by a distance $L$ and sinks by a depth $\delta$ when it stops. [(a), (c)] The size ratio between the sphere and the glass beads is not to scale.
@@
$\alpha=-5^{\circ}$ (-0.087 rad) 0.32-0.55 0.32-0.42 0.35-0.40 0.35-0.45
$\alpha=-10^{\circ}$ (-0.17 rad) 0.45-0.7 0.45-0.78 0.56-0.9 0.6-0.95
$\alpha=-15^{\circ}$ (-0.26 rad) 0.37-0.60 0.36-0.67 0.46-0.7 0.6-0.85
$\alpha=-20^{\circ}$ (-0.35 rad) 0.35-0.5 0.43-0.6 0.44-0.7 0.5-0.77
@@ 3 Analysis
The typical examples of actual movies are provided as supplementary material. As shown in Fig. 2 (a), a hemisphere of each sphere, except for the glass sphere, is painted red to detect the rolling orientation. Due to the surface condition of the glass sphere, it is not possible to paint it red with a permanent marker. The total rotational angle during $0\leq x\leq L$ is defined as $\theta_{\mathrm{stop}}$ (including the number of full rotations (each corresponding to $2\pi$)) and is measured in the counterclockwise direction by detecting the slope of the boundary between the white and red regions except for the glass sphere. To simplify the analysis, we evaluate only the initial and final images, together with the total number of rotations, which are manually counted. Only from the initial and final angles and rotation number, we calculate the total rotational angle $\theta_{\mathrm{stop}}$. Note that we do not measure the time-resolved rotational angle. By detecting the center of the sphere in each image, as shown in Fig. 2 (b), we obtain the sinking depth $\delta$ in the direction of gravity and the translational trajectory of the sphere, $X(t)$, where $t$ is the elapsed time after entering the granular slope.
@@ 4 Result
The measured sinking depth normalized by the sphere radius, $\delta/R$, is plotted in Fig. 3. As shown in Fig. 3 (a), $\delta/R$ is almost independent of $v_{0}$. There is an increasing trend beyond the error bars at $\alpha=-15^{\circ}$, but otherwise it is almost constant within the error bars regardless of initial velocity $v_{0}$. Therefore, in Fig. 3 (b)–(e), the average is taken over data with different $v_{0}$, and it can be seen that it does not depend on $\alpha$, similar to $v_{0}$. As an overall trend, the average value of $\delta/R$ depends on the density of the sphere, $\rho_{\mathrm{s}}$. As shown in Fig. 4, $\delta/R$ and $\rho_{\mathrm{s}}/\rho_{\mathrm{g}}$ exhibit the positive correlation. The density dependence of sinking depth has been investigated previously [29, 6, 10]. For a sphere dropped onto a granular surface without initial velocity, the normalized sinking depth was found to scale with the density ratio $(\rho_{\mathrm{s}}/\rho_{\mathrm{g}})$ as obtained by Uehara et al. [29],
@@
where $C_{\rho}$ is a dimensionless constant that depends on the friction coefficient between grains. Uehara et al. reported $C_{\rho}=0.51$ for a flat horizontal glass-bead surface [29], while Texier et al. obtained $C_{\rho}=0.61$ for a granular slope [6]. Although the value of $C_{\rho}$ differs, this scaling law [Eq. (1)] can also be applied to the sinking depth after a sphere rolled up a granular slope ($C_{\rho}=0.46\pm 0.05$[10]. When assuming that this scaling relation [Eq. (1)] is also applicable to the present study, we find that the relation is satisfied with $C_{\rho}=0.38\pm 0.06$, as shown in Fig. 4. Therefore, this scaling law is valid for spheres rolling on granular surfaces, and also the coefficient $C_{\rho}$ does not vary significantly between the rolling-down and rolling-up cases [10].
@@
Figure 4: The double logarithmic plot of $\delta/R$ vs $\rho_{\mathrm{s}}/\rho_{\mathrm{g}}$ for both downhill and uphill. We use the data (uphill) obtained by Fukumoto et al. [10]. Error bars indicate the standard deviation of various $\alpha,v_{0}$ conditions. The blue and red dashed line indicate the scaling relation [Eq. (1)] with $C_{\mathrm{\rho}}=0.38$ (R=0.78) and 0.46 (R=0.86), respectively. Here, R is the correlation coefficient.
@@
The slip ratio $s$ is computed by substituting the measured values of $L$ and $\theta_{\mathrm{stop}}$ into Eq. (2). Note that when the sphere is rolling down a slope, the translational and rotational motions stop almost simultaneously. As shown in Fig. 7, $s$ increases with $\alpha$, where the definition of $\pm\alpha$ corresponds to rolling up $(+)$ and rolling down ()(-). (For rolling up $(+)$, we use the data obtained by Fukumoto et al. [10].) The $s$ and $\alpha$ show the positive correlation as an overall trend. Moreover, the variation of $s$ appears to be roughly continuous for each sphere as $\alpha$ changes. The mechanism governing slipping behavior is complex. The shear imposed by the sphere rolling develops a localized shear zone called shear banding structure. This related phenomenon is reported by [19] The relation between shear banding structure and slipping behavior should be revealed to understand the fundamental physical process of energy dissipation. However, shear strain field in the glass beads bed cannot be measured in this experimental setup. Clarifying this mechanism remains an important issue for future studies.
As shown in Fig. 8, a bump is formed in front of the sphere, and we will focus on the relationship between this bump and slip ratio. Even though the sinking depth changes little between rolling up and rolling down, the deformation of the granular surface in front of the sphere is larger when it rolls down a granular slope, as shown in Fig. 8. During rolling down motion, the sphere is supposed to accelerate under gravity but the sphere actually decelerates due to the effect of the bump of glass beads accumulating in front of it. The bumpy structure produces a resistive force opposing its motion. In contrast, during rolling up motion, the deceleration due to gravity dominates over the resistance from the granular bed. These observations indicate that the size of bump is larger during rolling down motion than during rolling up motion. Numerical calculations have shown that the larger the slip ratio of a wheel rolling on granular ground, the smaller the bump formation (terrain deformation formed in front of the wheel) [15]. To purely evaluate the effect of $s$ on this phenomenon, data from different materials could be used. However, material density also affects the sinking depth. To examine the effect of bump formation under identical sinking depths, it is more appropriate to compare data obtained using the same material but different $\alpha$ values. Considering these points, the above-discussed tendency between bump formation and slip ratio is consistent with our observations. To quantitatively discuss this bump formation effect, much more precise measurements of granular surface deformation are required. This is one of the most important issues for future study.
Figure 7: The relation between slip ratio $s$ and angle $\alpha$. $\alpha<0$ and $\alpha>0$ correspond to rolling down and up cases respectively. We use the data ($\alpha>0$) obtained by Fukumoto et al. [10]. Error bars indicate the standard deviation of all the data. R=0.93(polyethylene), 0.97(polyacetal), 0.95(alumina ceramic), R: the correlation coefficient.